OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____.

OpenStudy (anonymous):

The top equation multiplied by 2 = 4x+2y=6. Add that to the bottom one and you get: 5x=5. So, x=1, then sub that back into the top or bottom (I'll do top) and you get 2+y=3. Therefore, y=1. Check these values by plugging them back into both equations and you see they are correct!

OpenStudy (anonymous):

) (x + y = 4) x 2 = 2x + 2y = 8 2x + 2y + x - 2y = 8 - 5 ( adding the above equation with the 2nd given equation ) 3x = 3 2) 2x + y = -1 X - y = -5 Multiply the 2nd equation by 2. 2x - 2y = -10 Subtract this equation from the 1st given equation. 2x + y - 2x + 2y = -1 +10 3y = 9 Therefore, y = 9 / 3 = 3 Substitute y = 3 in the 2nd given equation. X - 3 = -5 Therefore, x = -5 +3 = -2 Therefore, the answer is ( -2, 3 ). 3) -2u^3v +5uv -1 +7u^3v -2u^2v^2 +5u^2v^2 -uv +6 = = ( -2u^3v +7u^3v ) + ( 5uv -uv ) + ( -2u^2v^2 + 5u^2v^2 ) + ( -1 +6 ) = 5u^3v + 3u^2v^2 + 4uv + 5.